Identify the vertex,the focus ,directrix ,axis of symmetry, and the latus rectum
1.x=1/24y^2
2.y=-1/4x^2

Answer by lwsshak3(11628)

Identify the vertex,the focus ,directrix ,axis of symmetry, and the latus rectum
1.x=1/24y^2
y^2=24x
parabola opens right:
basic form of equation:(y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex
axis of symmetry:y=0, or x-axis
vertex:(0,0)
4p=24
p=6
focus: (6,0)
directrix: x=-6
latus rectum or focal width=4p=24
..
2.y=-1/4x^2
x^2=-4y
parabola opens down:
basic form of equation:(x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex
axis of symmetry:x=0, or y-axis
vertex:(0,0)
4p=4
p=1
focus: (0,1)
directrix: x=-1
latus rectum or focal width=4p=4